Which of these sentences are propositions? What are the truth values of those that are propositions?


1
Expert's answer
2021-10-12T09:41:20-0400

QUESTION

Which of these sentences are propositions? What are the truth values of those that are propositions?a) Boston is the capital of Massachusetts.b) Miami is the capital of Florida.c) 2 + 3 = 5.d) 5 + 7 = 10.e) x + 2 = 11.f ) Answer this question.

SOLUTION

Step-by-step explanation:

Proposition:

  • It is a declarative statement that is either false or true.
  • It cannot be both true or false.

a) Boston is the capital of Massachusetts.

The given statement is true. Hence, the given statement is a proposition.

"ANSWER = True"

b) Miami is the capital of Florida.

The given statement is false. Hence, the given statement is a proposition.

"ANSWER = False"

c) "2 + 3 = 5."

The given statement is true. Hence, the given statement is a proposition.

"ANSWER = True"

d) "5 + 7 = 10."

The given statement is false. Hence, the given statement is a proposition.

"ANSWER = False"

e) "x + 2 = 11."

The given statement can neither be true or false. It depends on the value of x. Hence, it is not a proposition.

"ANSWER =" Not a proposition

f ) Answer this question.

The given statement is not a declarative in nature. Hence, it is not a proposition.

"ANSWER =" Not a proposition